OpenStudy (anonymous):

Simplify 4 x all over 9 plus quantity x minus 10 over 9 (1 point)

OpenStudy (asnaseer):

elyise: Welcome to OpenStudy! Since this is your first question I assume you are not familiar with the site yet. The idea is to post one question at a time and at least show some attempt at solving the question so that others can spot where you may be having difficulties and guide you to the solution. Thank you.

OpenStudy (anonymous):

Okay thanks for the tips!

OpenStudy (asnaseer):

yw :)

OpenStudy (asnaseer):

I can help you with the first one - but please post the others (one at a time) as separate questions after this. You should be able to edit your question and remove al but the first question from here.

OpenStudy (asnaseer):

so the first question is to simplify this - correct?\[\frac{4x}{9}-\frac{x-10}{9}\]

OpenStudy (anonymous):

correct

OpenStudy (asnaseer):

ok, first thing to notice is that both fractions have the same denominator (9). this means that you can combine the numerators above it as follows:\[\frac{4x}{9}-\frac{x-10}{9}=\frac{4x-(x-10)}{9}\]understand?

OpenStudy (anonymous):

makes sence

OpenStudy (asnaseer):

ok, now see if you can simplify the numerator by removing the braces

OpenStudy (anonymous):

4x^2-10/9?

OpenStudy (asnaseer):

you have multiplied here which is not correct. the first step is to just remove the braces from:\[4x-(x-10)\]this then becomes:\[4x-(x-10)=4x-x+10\]

OpenStudy (anonymous):

ooooh okay that makes more sence lol

OpenStudy (asnaseer):

the -10 inside the braces becomes +10 because of the rule (minus and minus makes a plus)

OpenStudy (anonymous):

okay then what?

OpenStudy (asnaseer):

now combine "like" terms in:\[4x-x+10\]what do you get?

OpenStudy (anonymous):

3x+10?

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

so is that my answer??

OpenStudy (asnaseer):

so we end up with:\[\frac{4x}{9}-\frac{x-10}{9}=\frac{4x-(x-10)}{9}=\frac{4x-x+10}{9}=\frac{3x+10}{9}\]

OpenStudy (anonymous):

ah okay.... i was wondering what we did with the 9 haha well thanks!! :)

OpenStudy (asnaseer):

yw :)